# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v2

@@ 4.3 Graph Regression
We experiment with two molecular datasets: (1) ZINC-12K [18], and (2) QM9 [32]. ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important chemical property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen (H), Carbon (C), Oxygen (O), Nitrogen (N), and Flourine (F) atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation and a test set according to a $80\%/10\%/10\%$ split.
@@
We perform grid search to select values for the different hyperparameters. For both models, we choose the number of hidden layers from $\{2,3,4\}$, and the learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [19]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer which maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
Table 5: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on the ZINC-12K and QM9 datasets.
@@ Results.
The results are shown in Table 5. We observe that on both considered datasets, KAGIN significantly outperforms the GIN model. Note that these datasets are significantly larger (in terms of number of samples) compared to the graph classification datasets of Table 3. More specifically, KAGIN offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN. Those improvements suggest that KANs might be more effective than MLPs in regression tasks.
@@ 5 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. In the classification tasks, there does not appear to be a clear winner, with both architectures outperforming each other on some datasets. In the graph regression task, however, preliminary results seem to indicate that KAN has an advantage over MLP. This paper shows, through its preliminary results, that such KAN-based GNNs are valid alternatives to the traditional MLP-based models. We thus believe that these models deserve the attention of the graph machine-learning community.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove highly relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Finally, a thorough study of the effect of the different hyperparameters could be leveraged, allowing to fully exploit the richness of splines while retaining small networks.
6 Acknowledgements
This work was partially supported by the The Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and partner Swedish universities and industry.

### v3

@@ 4.3 Graph Regression
We experiment with two molecular datasets: (1) ZINC-12K [42], and (2) QM9 [43]. ZINC-12K consists of $12,000$ molecules. The task is to predict the constrained solubility of molecules, an important property for designing generative GNNs for molecules. The dataset is already split into training, validation and test sets ($10,000$, $1,000$ and $1,000$ graphs in the training, validation and test sets, respectively). QM9 contains approximately $134,000$ organic molecules. Each molecule consists of Hydrogen, Carbon, Oxygen, Nitrogen, and Flourine atoms and contain up to 9 heavy (non Hydrogen) atoms. The task is to predict $12$ target properties for each molecule. The dataset was divided into a training, a validation, and a test set according to a $80\%/10\%/10\%$ split.
@@
We perform grid search to select values for the different hyperparameters. We choose the number of hidden layers from $\{2,3,4\}$ for GIN and $\{1,2,3\}$ for KAN. We use different ranges for each model was because we wanted to limit the size advantage of KAN against MLP, since a layer with the same width of KAN has more parameters than an MLP one). For both models we search a learning rate from $\{10^{-3},10^{-4}\}$. For GIN, we choose the hidden dimension size from $\{32,64,128,256,512,1024\}$, while for KAGIN, we choose it from $\{4,8,16,32,64,128,256\}$. For KAGIN, we also select the grid size from $\{1,3,5,8,10\}$ and for BS-KAGIN the spline order from $\{3,5\}$. To produce graph representations, we use the sum operator. The emerging graph representations are finally fed to a 2-layer MLP (for GIN) or a 2-layer KAN (for KAGIN) which produces the output. We set the batch size equal to $128$ for all models. We train each model for $1,000$ epochs by minimizing the mean absolute error (MAE). We use the Adam optimizer for model training [39]. We also use early stopping with a patience of $20$ epochs. For ZINC-12K, we also use an embedding layer that maps node features into $100$-dimensional vectors. We choose the configuration that achieves the lowest validation error. Once the best configuration is found, we run $10$ experiments and report the average performance on the test set. For both datasets and models, we set the number of message-passing layers to $4$. On QM9, we performed a joint regression of the $12$ targets.
Table 6: Average MAE ($\pm$ standard deviation) of the KAGIN and GIN models on graph regression.
@@ Results.
The results are shown in Table 6. We observe that on both considered datasets, both KAGIN variants significantly outperform the GIN model. Note that these datasets are significantly larger (in terms of the number of samples) compared to the graph classification datasets of Table 5. The results also indicate that BS-KAGIN outperforms RBF-KAGIN on both ZINC-12K and QM9. However, the difference in performance between the two models is very small. Thus, it appears that both B-splines and RBFs are good candidates as KANs’ base components in regression tasks. In addition, we should note that the BS-KAGIN model offers an absolute improvement of approximately $0.11$ and $0.03$ in MAE over GIN. Those improvements suggest that KANs might be more effective than MLPs in regression tasks.
5 Analysis of Size and Time
5.1 Selected models
Following the above results, and it is legitimate to wonder whether the performance of the KAN models comes from a larger size than their MLP-based counterparts. We show in this section that it is not the case. Using the graph-classification models trained for the results provided in 4.2, we display here three relations: the size-to-performance relation, the size-to-training time relation and the training time-to-performance relation.
Figures 2 and 4 show these relations for datasets NCI1 and IMDB-MULTI respectively. We also show in 8 and 8 the numerical values associated to these graphs. The rest of the graph classification datasets are provided in App. 5.2. A first thing that we can see from both figures is that, for the same parameter count, B-splines-based KANs are significantly slower than both other architectures. RBF-based KANs are much faster, showing that this implementation succeeds in alleviating the computational bottleneck induced by the sequential computation necessary for B-splines. Finally, MLPs are the fastest for equal parameter counts. While these results were expected, we notice in both the tables and figures that RBF-KAGINmodels end up significantly smaller than their MLP-based counterparts, while maintaining equal or better performance, and about equal training time. On the other hand, the B-Splines based method is significantly slower, by sometimes a 5 fold factor. This establishes the RBF-based architecture as the best compromise between scalability and performance, although further work could be made in studying the performance of very small KAN-based against MLP-based models.
Figure 1: Parameters count/size/performance comparison, GIN, NCI1
Figure 2: Parameters count/size/performance comparison, GCN, NCI1
Figure 3: Parameters count/size/performance comparison, GIN, IMDB-MULTI
Figure 4: Parameters count/size/performance comparison, GCN, IMDB-MULTI
Table 7: Time and parameter sizes for selected models, NCI1
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.794 $\pm$ 0.016 572338.4 2.220
RBF-KAGIN 0.783 $\pm$ 0.017 139168.0 0.358
GIN 0.783 $\pm$ 0.024 1337369.8 0.359
BS-KAGCN 0.716 $\pm$ 0.031 90196.5 0.752
RBF-KAGCN 0.640 $\pm$ 0.075 166960.5 0.356
GCN 0.604 $\pm$ 0.046 379634.0 0.277
Table 8: Time and parameter sizes for selected models, IMDB-MULTI
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.499 $\pm$ 0.033 296418.8 0.367
RBF-KAGIN 0.488 $\pm$ 0.043 48705.1 0.169
GCN 0.503 $\pm$ 0.038 330987.6 0.146
BS-KAGIN 0.500 $\pm$ 0.042 46052.3 0.216
RBF-KAGIN 0.485 $\pm$ 0.038 40118.5 0.161
MLP 0.462 $\pm$ 0.046 111911.4 0.143
5.2 Theoretical models
We give in Table 9 the training time per epoch for different configurations of the BS-KAGIN, RBF-KAGIN and GIN models on a graph-level task. We observe that for a given number of message-passing layers and hidden dimension size, GIN is the fastest, followed by RBF-KAGIN. BS-KAGIN is noticeably slower than the others. It suffers doubly: its parameter count increases much quicker than both other models for same hyperparameters; but even for similar numbers of parameters, BS-KAGIN is by far the slowest. Spline order seems to have a larger impact on running time than grid size. This shows that RBFs do alleviate some of the computational bottlenecks of the B-splines from the original KAN paper. Table 16 in Appendix C shows similar observations on a node-level task.
Table 9: Training time per epoch on NCI1 dataset. $5$ message-passing layers. batch size set to $128$.
Architecture
Hidden
Dimension
Hidden
Layers
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 256 2 NA NA 670722 0.303
GIN 256 4 NA NA 1465346 0.390
GIN 512 2 NA NA 2652162 0.541
GIN 512 4 NA NA 5814274 0.951
RBF-KAGIN 256 1 1 NA 822872 0.386
RBF-KAGIN 256 1 4 NA 1639274 0.599
RBF-KAGIN 256 3 1 NA 3191408 0.783
RBF-KAGIN 256 3 4 NA 6367142 1.577
RBF-KAGIN 512 1 1 NA 3218520 0.807
RBF-KAGIN 512 1 4 NA 6424170 1.441
RBF-KAGIN 512 3 1 NA 12674160 2.198
RBF-KAGIN 512 3 4 NA 25317030 4.429
BS-KAGIN 256 1 1 1 1091072 0.446
BS-KAGIN 256 1 1 4 1907456 2.964
BS-KAGIN 256 1 4 1 1907456 0.720
BS-KAGIN 256 1 4 4 2723840 4.240
BS-KAGIN 256 3 1 1 4236800 0.937
BS-KAGIN 256 3 1 4 7412480 9.703
BS-KAGIN 256 3 4 1 7412480 1.963
BS-KAGIN 256 3 4 4 10588160 14.572
6 Discussion
Prominent application domains.
KANs, both in our work and in the recent literature, seem to be a promising alternative to MLPs for regression tasks. They rely on combinations of continuous functions which could explain their improved performance for continuous outputs. Moreover, this property and their empirical performance in regression tasks renders them an ideal candidate for physics-inspired architectures [44], which address problems of high importance, from climate forecasting [45] to atomic simulations [46] and rely heavily on graph-based architectures.
Explainability matters.
The potential of explainability for KAN-based models is arguably better compared to that of MLPs with post-hoc methods (e.g. Shapley values, gradient-based), since closed-form symbolic formulas can be drawn from the former, providing complete overview of the underlying computations. Hence, KAN-based graph models could potentially prove more friendly to practitioners from application fields, such as biology and physics, who tend to be skeptical of the black-box approach.
GNNs for fixed graphs inspired from KANs.
Each KAN architecture consists of a set of neurons. Each one of those neurons is connected to other neurons. Thus, KANs can be naturally modeled as graphs where nodes represent neurons and edges represent the connections between them. Each layer of a KAN is actually a complete bipartite graph, while a KAN has learnable activation functions on the edges of its associated graph. There actually exist datasets that contain multiple graphs which share the same graph structure, but differ in the node and edge features. Two examples of such datasets are the MNIST and CIFAR10 graph classification datasets [47]. Since the graph structure is fixed (and identical for all graphs), we could associate each edge of the graph with one or more learnable activation functions (shared across graphs), similar to what KANs do. This would allow the model to learn different functions for each edge of the graph, leading potentially to high-quality node and graph representations.
Advantages and future work.
Finally, we discuss potential advantages that KANs might have over MLPs, and leave their investigation for future work. First, their ability to accurately fit smooth functions could prove relevant on datasets where variables interact with some regular patterns. Second, their interpretability could be leveraged to provide explanations on learned models, giving insights into the nature of interactions among entities. Finally, a thorough study of the effect of the different hyperparameters could be pursued, allowing to fully exploit the splines or radial basis functions while retaining small networks.
@@ 7 Conclusion
In this paper, we have investigated the potential of Kolmogorov-Arnold networks in graph learning tasks. Since the KAN architecture is a natural alternative to the MLP, we developed two GNN architectures, KAGCN and KAGIN, respectively analogous to the GCN and GIN models. We then compared those architectures against each other in both node- and graph-level tasks. For graph and node classification, there does not appear to be a clear winner. For graph regression, our results indicate that KAN has an advantage over MLP, with the B-spline variant proving superior in terms of performance. Nonetheless, this gain comes at the cost of computational efficiency. In this regard, the RBF-based vairant offers a slightly lesser performance improvement, but proves much more efficient in terms of training time.
This paper shows that such KAN-based GNNs are valid alternatives to the traditional MLP-based models. We thus believe that these models deserve the attention of the graph machine-learning community, and that future work should be carried out to identify their specific advantages and weaknesses, allowing for an enlightened choice between KANs and MLPs based on the task and context at hand.
Appendix A Further Graph Classification Experiments
Besides the graph classification experiments that are presented in the main paper, we also experimented with the Peptides-func dataset, a graph classification dataset from the Long Range Graph Benchmark [48]. The dataset consists of $15,535$ graphs where the average number of nodes per graph is equal to $150.94$. We evaluated GIN, BS-KAGIN and RBF-KAGIN on this dataset. We performed a grid search where for all models we chose the learning rate from $\{0.001,0.0001\}$, the number of hidden layers of MLPs and KANs from $\{1,2,3\}$ and the hidden dimension from $\{64,128\}$. We set the number of neighborhood aggregation layers to $5$ for all models. For both BS-KAGIN and RBF-KAGIN, grid size was chosen from $\{3,5\}$, and for BS-KAGIN, spline order was chosen from $\{3,5\}$. Each experiment was run $4$ times with $4$ different seeds and Table 10 illustrates the average precision achieved by the three models.
Table 10: Average Precision ($\pm$ standard deviation) of the KAGIN and GIN models on the Peptides-func dataset.
Peptides-func
GIN 55.01 $\pm$ 0.75
BS-KAGIN
58.71
$\pm$ 1.02
RBF-KAGIN 57.68 $\pm$ 2.05
We observe that BS-KAGIN and RBF-KAGIN outperform GIN on the Peptides-func dataset. Both models offer an absolute improvement of more than $2.5\%$ in average precision over GIN. Thus, even in tasks where long-range interactions between nodes need to be captured, for a fixed receptive field, KAN layers can help the models better capture those interactions.
Appendix B Time/Performance/Accuracy analysis
Figures 6 to 14 and tables 11 to 15 regarding the time/size/accuracy analysis on the graph classification task.
Table 11: Time and parameter sizes for selected models, IMDB-BINARY
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.734 $\pm$ 0.038 148236.3 0.225
RBF-KAGIN 0.720 $\pm$ 0.036 168383.6 0.129
GIN 0.731 $\pm$ 0.038 1172520.5 0.110
BS-KAGCN 0.708 $\pm$ 0.037 39053.5 0.150
RBF-KAGCN 0.726 $\pm$ 0.044 37322.5 0.109
GCN 0.687 $\pm$ 0.034 118826.3 0.099
Table 12: Time and parameter sizes for selected models, ENZYMES
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.427 $\pm$ 0.085 450136.6 0.336
RBF-KAGIN 0.412 $\pm$ 0.079 115346.8 0.061
GIN 0.533 $\pm$ 0.117 1623033.2 0.048
BS-KAGCN 0.577 $\pm$ 0.076 83636.3 0.103
RBF-KAGCN 0.594 $\pm$ 0.053 125854.4 0.050
GCN 0.221 $\pm$ 0.053 280040.0 0.039
Table 13: Time and parameter sizes for selected models, DD
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.754 $\pm$ 0.032 169417.4 1.206
RBF-KAGIN 0.757 $\pm$ 0.029 147005.1 0.236
GIN 0.748 $\pm$ 0.044 168957.4 0.144
BS-KAGCN 0.733 $\pm$ 0.042 110278.6 0.596
RBF-KAGCN 0.683 $\pm$ 0.043 110899.2 0.185
GCN 0.632 $\pm$ 0.036 342023.2 0.225
Table 14: Time and parameter sizes for selected models, MUTAG
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.851 $\pm$ 0.058 236235.8 0.056
RBF-KAGIN 0.855 $\pm$ 0.051 62530.0 0.020
GIN 0.853 $\pm$ 0.069 848213.0 0.014
BS-KAGCN 0.755 $\pm$ 0.077 19496.1 0.022
RBF-KAGCN 0.735 $\pm$ 0.110 47769.4 0.017
GCN 0.563 $\pm$ 0.143 72072.2 0.012
Table 15: Time and parameter sizes for selected models, PROTEINS_full
Model Accuracy Parameter count time per epoch (s)
BS-KAGIN 0.741 $\pm$ 0.034 321367.7 0.472
RBF-KAGIN 0.733 $\pm$ 0.034 51498.8 0.096
MLP 0.710 $\pm$ 0.027 942781.0 0.073
BS-KAGCN 0.756 $\pm$ 0.034 37688.7 0.133
RBF-KAGCN 0.742 $\pm$ 0.035 48950.4 0.078
GCN 0.605 $\pm$ 0.020 8199.8 0.061
Figure 5: Parameters count/size/performance comparison, GIN, IMDB-BINARY
Figure 6: Parameters count/size/performance comparison, GCN, IMDB-BINARY
Figure 7: Parameters count/size/performance comparison, GIN, ENZYMES
Figure 8: Parameters count/size/performance comparison, GCN, ENZYMES
Figure 9: Parameters count/size/performance comparison, GIN, DD
Figure 10: Parameters count/size/performance comparison, GCN, DD
Figure 11: Parameters count/size/performance comparison, GIN, MUTAG
Figure 12: Parameters count/size/performance comparison, GCN, MUTAG
Figure 13: Parameters count/size/performance comparison, GIN, PROTEINS_full
Figure 14: Parameters count/size/performance comparison, GCN, PROTEINS_full
Appendix C Training Times - Node Classification
Table 16: Training time per epoch on the node-classification task (Cora dataset, 3 MP-layers).
Architecture
Hidden
Dimension
Hidden
Layers
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GCN 256 NA NA NA 378934 0.003
GCN 1024 NA NA NA 1485622 0.005
BS-KAGCN 256 NA 1 1 1514684 0.017
BS-KAGCN 256 NA 1 4 2650697 0.160
BS-KAGCN 256 NA 4 1 2650697 0.031
BS-KAGCN 256 NA 4 4 3786710 0.222
BS-KAGCN 512 NA 1 1 2989244 0.020
BS-KAGCN 512 NA 1 4 5231177 0.179
BS-KAGCN 512 NA 4 1 5231177 0.039
BS-KAGCN 512 NA 4 4 7473110 0.254
RBF-KAGCN 256 NA 1 NA 1142524 0.016
RBF-KAGCN 256 NA 4 NA 2278543 0.031
RBF-KAGCN 1024 NA 1 NA 4462588 0.029
RBF-KAGCN 1024 NA 4 NA 8916367 0.057
GIN 256 2 NA NA 867335 0.007
GIN 256 4 NA NA 1130503 0.008
GIN 1024 2 NA NA 5042183 0.020
GIN 1024 4 NA NA 9240583 0.033
BS-KAGIN 256 1 1 1 1514684 0.017
BS-KAGIN 256 1 1 4 2650697 0.158
BS-KAGIN 256 1 4 1 2650697 0.032
BS-KAGIN 256 1 4 4 3786710 0.224
BS-KAGIN 256 3 1 1 3990528 0.025
BS-KAGIN 256 3 1 4 6983424 0.229
BS-KAGIN 256 3 4 1 6983424 0.051
BS-KAGIN 256 3 4 4 9976320 0.334
BS-KAGIN 512 1 1 1 2989244 0.021
BS-KAGIN 512 1 1 4 5231177 0.184
BS-KAGIN 512 1 4 1 5231177 0.041
BS-KAGIN 512 1 4 4 7473110 0.261
BS-KAGIN 512 3 1 1 10078208 0.046
BS-KAGIN 512 3 1 4 17636864 0.354
BS-KAGIN 512 3 4 1 17636864 0.089
BS-KAGIN 512 3 4 4 25195520 0.499
RBF-KAGIN 256 1 1 NA 1142524 0.017
RBF-KAGIN 256 1 4 NA 2278543 0.032
RBF-KAGIN 256 3 1 NA 3002487 0.023
RBF-KAGIN 256 3 4 NA 5995401 0.046
RBF-KAGIN 1024 1 1 NA 4462588 0.03
RBF-KAGIN 1024 1 4 NA 8916367 0.057
RBF-KAGIN 1024 3 1 NA 21429879 0.091
RBF-KAGIN 1024 3 4 NA 42838665 0.185